Consider two lines L1 and L2 whose direction cosines l1,m1,n1 and l2,m2,n2 are given by the equations for l,m,n :
al +bm+cn = 0, f mn+gnl +hlm = 0,
where abc not = 0. Show that if L1 ⊥ L2, then f/a +g/b +h/c = 0.
1
Expert's answer
2020-05-18T19:13:58-0400

We have given that "L_1 \& L_2" are equation of line whose direction cosines are given as "(l_1,m_1,n_1) \& (l_2,m_2,n_2)" respectively obtained from solution of two given equation as follows,


"al +bm+cn = 0 \hspace{1cm} (1)\\ f mn+gnl +hlm = 0 \hspace{1cm}(2)"

where "abc \neq 0" ,also given that "L_1 \bot L_2 \implies l_1l_2+m_1m_2+n_1n_2=0 \: (\star)" .

Let's eliminate one of the variable from "(1)" ,say "l" ,thus "l=-\frac{bm+cn}{a}" and substitute to "(2)" we get,


"fmn+gn(-\frac{bm+cn}{a})+hm(-\frac{bm+cn}{a})=0\\
\implies hbm^2+(hc+gb-af)mn+gcn^2=0\\
\implies hby^2+(hc+gb-af)y+gc=0 \hspace{1cm}(\clubs)"


where

"y=\frac{m}{n}"



Clearly, equation "(\clubs)" is quadratic in "y" ,hence by relation between roots and coefficients(Vieta's Theorem) we get,


"y_1+y_2=-\frac{(hc+gb-af)}{hb} \hspace{1cm} (3)\\
y_1y_2=\frac{gc}{hb}\hspace{1cm} (4)"

Hence,from equation "(4)" ,

"\frac{m_1}{n_1}\frac{m_2}{n_2}=\frac{gc}{hb}\\
\implies \frac{m_1m_2}{\frac{g}{b}}=\frac{n_1n_2}{\frac{h}{c}} \hspace{1cm}(5)"

In the same manner if we eliminate other two variables "n ,m" (say "n" ) and do the same calculation as above we get,

"\frac{m_1m_2}{\frac{g}{b}}=\frac{l_1l_2}{\frac{f}{a}} \hspace{1cm}(6)"

Now from equation "5" and "6" we finally get


"\frac{m_1m_2}{\frac{g}{b}}=\frac{l_1l_2}{\frac{f}{a}} =\frac{n_1n_2}{\frac{h}{c}} \hspace{1cm}(7)"

Thus from equation "7" if we consider "\frac{m_1m_2}{\frac{g}{b}}=\frac{l_1l_2}{\frac{f}{a}} =\frac{n_1n_2}{\frac{h}{c}}=p" ,clearly "p \neq 0" ,hence,


"l_1l_2+m_1m_2+n_1n_2=p\bigg(\frac{g}{b}+\frac{f}{a}+\frac{h}{c}\bigg )"

but from equation "(\star)" we get,


"l_1l_2+m_1m_2+n_1n_2=p\bigg(\frac{g}{b}+\frac{f}{a}+\frac{h}{c}\bigg )=0\\
\implies p\bigg(\frac{g}{b}+\frac{f}{a}+\frac{h}{c}\bigg )=0\\
\implies \bigg(\frac{g}{b}+\frac{f}{a}+\frac{h}{c}\bigg )=0 \hspace{1cm}(\because p\neq0)\\"

Therefore we are done.